Nickblade10:

CAN ANYONE HELP ME?: Kate and Stella both worked at the coffee shop today. Kate's total cups of coffee made is represented by f(x); and Stella's total cups of coffee made is represented by g(x). Use the functions below to write a function that represents the total cups of coffee they made today. f(x) = 6x − 8 g(x) = 3x + 1 9x + 9 9x − 7 3x − 7 3x − 9

Extrinix:

Refer back to your previous post. https://questioncove.com/study#/updates/6391e5175606e1cd58e4c000

Nickblade10:

I know but everytime i do it, i get a answer thats not on the choices

Extrinix:

Extracted from the previous post:

@extrinix wrote:
So, normally speaking, if you're trying to get the total of two equations, you add everything together. \(\sf{f(x) = 6x − 8 \\ g(x) = 3x + 1}\) Given these, you're trying to do \(\sf{f(x)+g(x)}\) to get what your total would be (let's use \(\sf{t(x)}\) for the equaled). \(\sf{t(x)=(6x+3x) + ((-8)+1)}\) Once you solve those, you'll have the answer you need. Best of luck, \(\style{font Family:Arial; text-shadow: 2px 2px 3px white, 0 0 10px white, 0 0 5px white;color:#23B7F3;font-size:100px;padding:50px;width:500px;height:100%;position:absolute;left:0;z-index:500;text-align:center;} {\Huge\sf\color {} {{-~Akuma~matata}}}\) "what a wonderful phrase"

Extrinix:

Here, you need to calculate this. \(\sf{t(x)=(6x+3x) + ((-8)+1)}\) Try doing it like this. \(\sf{~~~6+3 = y}\) \(\sf{-8+1=z}\) and then just plug it back into the equation, with it's combined value \(\sf{t(x)=(y)x + (z)}\)

Nickblade10:

how do you get the combined vaule?

Noobsabotes:

So this is the answer cause if it is I make sure I work with these problems

Nickblade10:

So, to let you all know this answer is 9x-7, supplied by a good friend

Extrinix:

I can't give the answer, you have to figure that out yourself. It's simple math, you just have to take the time and effort to work it out.

Extrinix:

My bad, bad wifi

Nickblade10:

i did, you have no idea how bad ar math i am, and i really appreciate your help, its why i gave you another medal